INAF LBC+LUCIFER queue observing, 2010 February13-14

Observer: V. Testa, K.Boutsia
Telescope Operator: WWack
Support Astronomer: D. THompson
Instrument: LBC

Summary

* Good weather, thin clouds at the horizon.
* El. axis problem soon after opening. Evening flat fields lost
* Time lost: 90 min telescope
* Done standard star middle and end of night, the last during civil twilight

Data Summary

Project Objects Filter ExpTime NotesSorted descending
Calibration SA 107 Us,U,B,g,V,r 0.6h two airmasses
Calibration SA104 Us,U,B,g 0.4h two airmasses
Giallongo Q0933 g 3.3h aver. seeing 1.0"
Rigliaco NGC 2264 B, U 0.8 h aver. seeing 0.9"
Giallongo Cosmos g 0.5h  
Mereghetti LDTR Uspec 0.5h  
Giallongo Q1623 Uspec 1.0h  

Details

01:05 - Open, try to point FF spot.

01:20 - El. axis fail to start

02:38 - Slew to Rigliaco field NGC 2264

02:39 - dofpia,/BlueOnly
  • -16752 0 0 0 0 0 0 2.33 6.60
  • exited on the second iteration. Strong coma.
  • clearly wrong. Put Z7=-2500, Z4=10000 and retry

02:46 - dofpia,/BlueOnly
  • Z4 still big but correctly fit. (+8900)
  • error again. Still residual coma. Try Z7 -1000
  • 1222 1234 -1837 -1891 -1393 -35 150 0.85 1.33
  • -1084 -1012 555 1196 -367 31 150 1.65 0.79
  • -701 454 139 -206 74 348 150 1.34 0.40
  • -839 -116 496 177 -122 273 122 1.67 0.42
  • -439 201 -161 48 -2 96 13 1.25 0.21

02:57 - Rigliaco_NGC2264_B1

  • 025906 - 0.87" seeing, round.
  • ~1.0 seeing at the end

03:37 - dofpia,/BlueOnly
  • -239 -153 535 689 7 -57 -150 0.94 0.37
  • 629 -321 -453 -90 241 -131 -98 1.02 0.34
  • -346 -196 -624 74 -10 184 -23 1.09 0.28
  • 511 -47 409 -87 18 -345 -89 0.98 0.33
  • 34 252 260 127 -73 -107 -5 0.99 0.16

03:45 - Rigliaco_NGC2264_U1

04:15 - Slew to Giallongo Q0933

04:17 - dopfia/BlueOnly
  • -3622 -10 99 261 -86 210 -24 1.03 1.44
  • 112 -117 -23 86 -75 -60 -71 1.09 0.12

04:22 - Q0933_g-SLOAN_giallongo_a

04:57 - dofpia/BlueOnly
  • 660 -42 297 -103 125 -336 -110 1.03 0.36
  • 277 37 -254 -415 -140 -86 -3 1.00 0.26
  • went out of synchro, not finding the image, and hence no pupils found

05:04 - ctrl/C backout Z11 and restart dofpia,/BlueOnly

  • no pupils found!!!!!

05:06 - dofpia,/BlueOnly,/X2
  • 248 -457 68 186 4 -247 -9 1.04 0.26
  • 97 186 -160 -36 -20 34 -50 0.95 0.12

05:10 - Q0933_g-SLOAN_giallongo_b

05:46 - dofpia,/BlueOnly,/X2
  • 865 262 -494 379 4 -345 39 1.18 0.48
  • -173 -377 865 6 -65 25 5 1.11 0.30
  • -161 -151 -336 -53 -23 66 27 1.23 0.16

05:51 - Q0933_g-SLOAN_giallongo_c Lost focus at the last iterations

06:27 - dofpia/Blueonly/X2
  • -411 216 3 -17 23 115 44 1.61 0.20

06:29 - Q0933_g-SLOAN_giallongo_d

  • Looses focus as OB goes by. Probably related to high elevation and failure of collimation model at -2.5C.
  • Quite strange because it did fine before.

07:04 - dofpia/BlueOnly/X2
  • -1427 444 100 14 346 258 48 1.27 0.63
  • 116 -523 79 151 326 14 4 1.35 0.24

07:09 - Q0933_g-SLOAN_giallongo_e

07:45 - dofpia,/BlueOnly,/X2
  • -545 -335 -415 200 -204 396 -41 1.02 0.40
  • 71 210 58 189 -10 -52 -79 1.01 0.14

07:49 - Q0933_g-SLOAN_giallongo_f

Always focus change during OB

08:25 - dofpia,/BlueOnly,/X2
  • -1432 -175 15 176 71 -205 16 1.09 0.59
  • -769 -171 -129 -196 -99 64 20 1.09 0.34
  • -439 471 42 -26 -322 87 -13 1.19 0.28

08:30 - Q0933_g-SLOAN_giallongo_g

Probably photometric. Do two standard fields, at low and high elevation

09:07 - SA107 UBVUsgr 4pos

09:29 - SA104 UBgUs 4 pos

09:41 - Slew to Cosmos field

09:43 - dofpia,/BlueOnly
  • -2760 -95 -220 373 -93 75 78 1.54 1.11
  • -1192 -247 130 75 7 251 -46 1.22 0.51
  • -125 -24 262 -15 98 6 -48 1.09 0.13

09:48 - cosmosG1_b

10:20 - Slew to Q1623 and collimate

10:22 - dofpia,/BlueOnly
  • -1431 564 139 74 13 -28 96 1.35 0.60
  • -323 70 239 -111 -163 48 38 1.32 0.19

10:26 - Q1623_Uspec_1

Low elevation yet. Seeing above 1.4", stop. Slew to Mereghetti

10:34 - slew to LDRT_Ui-1La

10:36 - dofpia,/BlueOnly
  • -831 -520 -568 159 143 131 16 1.25 0.43
  • -212 -60 -232 -1 -40 102 -78 0.98 0.15

10:39 - LDRT_Ui-1La (Mereghetti)
  • seeing ~1.2 , slightly above constraints improves during OB play
  • guiding star sometimes moves around

11:11 - Slew to Q1623 and dofpia,/BlueOnly
  • 2326 283 38 -63 -233 -83 14 1.12 0.93
  • 407 279 52 -52 -32 -59 13 1.08 0.20

11:14 - Q1623_Uspec_1

11:49 - dofpia,/BlueOnly
  • 1079 -275 383 -372 37 -284 29 1.12 0.52
  • -224 -188 425 145 100 167 4 1.08 0.22

11:53 - Q1623_Uspec_2
  • seeing 1.15 to 1.35, slightly above constraints

12:30 Slew to SA107 UBVUsgr 4pos and dofpia,/BlueOnly
  • -1997 11 -680 245 90 -7 83 0.96 0.82
  • -309 89 -159 -110 91 69 27 1.01 0.17

12:35 - SA107_UsUBgVr

12:51 - SA104_UBgUs

13:07 - Slew to Flat Field spot and wait the sky brightening

13:16 - SkyFlat_g 5x5s PA=0

13:20 - SkyFlat g 5x5s PA=180

Redo g 3s (~20000 cts)

13:32 - SkyFlat B 5x4s PA=0 last one slightly saturated

13:36 - SkyFlat B 5x2s PA=180 slightly saturated too....

13:41 - SkyFlat Uspec 5x1s PA=180

13:44 - Sky Flat U 5x0.7s PA=180

13:48 - Sky Flat Uspec 5x0.5s PA=0, saturates quite soon, aborted.

13:51 - End of Night

-- VincenzoTesta - 14 Feb 2010


Accidental overwrite with notes from Feb 14-15, can probably delete the following content (dthompson)

INAF LBC+LUCIFER queue observing, 2010 February14-15

Observer: V. Testa, K.Boutsia
Telescope Operator: WWack
Support Astronomer: D. Thompson
Instrument: LBC

Summary

* Fernando on the mountain. LBC Red pumped and LN2 filled. Temp. and pressure are decreasing regularly. * Good weather, very thin clouds at the horizon.

Data Summary

Project Objects Filter ExpTime Notes

Details

00:58 - 5 Bias monoB

01:08 - 10Bias_Bino

01:13 - stopped bias 8/10. Slew to 04:50 , +27:00 for flats

01:15 - Sky Flat Test U 0.3s (~62000)

01:18 - Sky Flat U 5x0.5s PA=0
  • 1st one slightly saturated

01:22 - Sky Flat Uspec 5x0.8s PA=0

01:25 - Sky Flat B 5x0.5s PA=0

01:29 - Sky Flat B 5x1.0s PA=180

01:33 - SkyFlat V 5x2.0s PA=0

01:38 - Sky Flat V 5x5.0s PA=180

01:43 - Sky Flat r 5x10.0s PA=180

  • 800 counts. Strange behavior, looks like shutter got stuck. Test with V and r fast OBs.
  • Shutter seems not to be working. Try shutdown and restarting the camera.

02:00 - No way. Shutter in strange position
  • Try to move with testpower and minicom.
  • Succeeded.

02:20 - Exposure ok. Restart (35 min downtime, although it does not affect science time, as it happened during twilight)

02:23 - Slew to NGC 2264

02:25 - dofpia,/BlueOnly
  • Usual Z7 -3000 Needed

02:28 - dofpia,/BlueOnly
  • Big pupils. Z7 not applied. Also need Z4 +10000

02:32 - dofpia,/BlueOnly
  • still not converging properly, sends crazy Z4=-13544
  • Z4 +14000
  • Z7 -2000

02:38 - try again....
  • reached a reasonable fit
  • 1785 -3111 -1172 1214 -970 -688 150 0.00 1.46
  • 414 1075 -473 -1262 -248 -184 150 1.93 0.71
  • 171 1201 -202 237 -224 41 150 0.89 0.40
  • -439 144 277 51 102 0 0 2.14 0.22

02:45 - NGC 2264 V1
  • Out of focus since the beginning.
  • Stop and redo dofpia

  • -154 278 281 308 81 448 103 1.07 0.35
  • -756 226 96 168 -85 248 37 1.28 0.36
  • -97 138 90 155 -7 45 -44 1.36 0.13

02:55 - NGC 2264 V1

03:35 - dofpia,/BlueOnly
  • -935 -117 -522 1021 -134 327 -150 1.07 0.65
  • 168 74 726 -93 114 -191 -141 0.99 0.28
  • 275 46 261 78 21 -75 -59 0.93 0.17

03:42 - NGC 2264 R1

-- VincenzoTesta - 15 Feb 2010
Topic revision: r1 - 31 Jul 2024, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback